OpenStudy (anonymous):

1. What is the electric force between a +2 µC point charge and a –2 µC point charge if they are separated by a distance of 5.0 cm? Show your work. (µC = 1.0 × 10–6 C)

OpenStudy (anonymous):

apply coulomb's law

OpenStudy (anonymous):

I know the formula for this, F=kq1q2/r^2 but I don't know how to apply it

OpenStudy (anonymous):

here q1 = 2uc and q2 = -2uc and r = 5*10^-2m

OpenStudy (anonymous):

What's k?

OpenStudy (anonymous):

k is a constant k = 8.98 *10^9

OpenStudy (anonymous):

So it would look like this? (8.98*10^9)(2)(-2)/5*10^-2

OpenStudy (irishboy123):

- micro coulombs - square the distance too , ie denominator

OpenStudy (shamim):

Given q1=2*10^-6coulomb q2=-2*10^-6 coulomb k=9*10^9N.m^2/C^2

OpenStudy (shamim):

F=?